OpenStudy (anonymous):

Solve the following equation: x divided by 8 equals twelve. x equals ninety-six x equals twenty x is greater than ninety-six x is less than ninety-six

OpenStudy (nightmarenight):

hOnk :0)

OpenStudy (anonymous):

\[\frac{ x }{ 8 }=12\] \[8\times \frac{ x }{ 8 }=12\times8\] x=12x8

OpenStudy (anonymous):

x=96

OpenStudy (anonymous):

Which one would it be then? c:

OpenStudy (anonymous):

x equals ninety-six